OpenStudy (anonymous):

The height of the tunnel at the center is 27 ft and the vertical clearance must be 9 ft at a point 24 ft from the center. Find an equation for the ellipse

OpenStudy (anonymous):

I got y^2 / a^2 + x^2 / b^2 = 1 y^2 / 729 + x^2 / b^2 = 1 81/729 + 576 / b^2 = 1 81/ 729 = 576/b^2 So what would the equation be?